OpenStudy (anonymous):

Multiply the quantity three point three times ten to the fourth power multiplied by the quantity three times ten to the eighteenth power. Express the answer in scientific notation.

OpenStudy (kropot72):

\[3.3\times 10^{4}\times 3\times 10^{18}=9.9\times 10^{22}\]